The Theatre Club sold a total of 415 tickets. The adult tickets cost $5 and the children's tickets cost $3. If $1,615 was collected, how many adult tickets were sold?
Answer by richwmiller(17219)

a+c=415,
5*a+3*c=1615
a=415-c
5*(415-c)+3*c=1615
2075-5c+3*c=1615
-2*c=-460
c=230
a=415-c
a=185 at $5 and c=230 at $3
check
185+230=415
5*a+3*c=1615
5*185+3*230=1615
925+690=1615
1615=1615
ok